OpenStudy (anonymous):

arrange them from highest to smallest .10.0 g, 10.0 cg, 1.00 kg, 1000.0 mg , 1000.0 dg

OpenStudy (redohawk):

okay, i can help i hope ^-^ gimme a sec :)

OpenStudy (anonymous):

okay .i will wait for your answer .i need it now .please?

OpenStudy (redohawk):

1) 1.00kg 2) 10.0g 3)1000.0mg 4)1000.0dg I don't know where cg would be..

OpenStudy (redohawk):

oh, in case i did it wrong. I recommand you take a look at this...

OpenStudy (redohawk):

Kg = Kilo gram = 1000g g = gram = 1g dg = decigram = 1/10 of a gram mg = milligram = 1/1000th of a g

OpenStudy (redohawk):

oh, and 1 cg = 0.01g

OpenStudy (anonymous):

it's ok with me if your answer is wrong .atleast you give your best .

OpenStudy (redohawk):

thanks for that. I just became your fan :D question for you? do you watch anime?

OpenStudy (anonymous):

yup.i love one piece.!!

OpenStudy (redohawk):

visit my channel, and maybe subscribe please ;D http://www.youtube.com/user/IzzyDoo1000

OpenStudy (anonymous):

oh.okay .i will .maybe sometimes . if it's ok?

OpenStudy (redohawk):

its awesome! I lof of poeple like it. I make amv's (anime music videos)

OpenStudy (anonymous):

hey .are u everyday OL?

OpenStudy (redohawk):

what is ol?

OpenStudy (anonymous):

OL .means online.

OpenStudy (redohawk):

ehh. i have not been on this for the past 7 months. but i am back now xD send me a message if you need help :)

OpenStudy (anonymous):

ahh. ok . thanks anyway.

OpenStudy (anonymous):

gud nyt .thanks again. :)

OpenStudy (mathstudent55):

Convert every amount to grams and compare them.

OpenStudy (anonymous):

Listen I won't give you the ans.. buy may lead you to it...

OpenStudy (mathstudent55):

A) \(10.0 ~g\) B) \(10.0 ~cg \times \dfrac{1~g}{100~cg} =0.1 ~g\) C) \(1.00 ~kg \times \dfrac{1000~g}{1~kg} = 1000~g\) D) \(1000.0 ~mg \times \dfrac{1~g}{1000~mg} = 1~g\) E) \(1000.0 ~dg \times \dfrac{1~g}{10~dg} = 100 ~g \)

OpenStudy (anonymous):

As mg is the samllest of all given units.. try converting all into mg... so you will only have to multiply.